1/2

Christian Morgenstern,  Christian Morgenstern,
The Banshee (An Approach)  El Hechicero
tradukita de Max Knight  tradukita de Hans Eiseneck

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por favor!”

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

“El hechiuno” le explica, “el hechidos” después indica, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

El hechicero, conmovido por lo que capta su oído, al maestro ruega, encantado, que le adjunte un quebrado.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. Hechiceros únicamente hay enteros.”

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

El brujo se volvió muy triste: “Mi cara mitad, ¿no la viste?” Ya que no era sabio con gratitud se despidió.

…  …

2/2

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

http://www.poezio.net/version?poem-id=825&version-id=1660,1656